Nongravitating vacuum energy and the volume elements of spacetime^{∗}

Frank Gronwald^{†}, Uwe Muench^{‡}, Friedrich W. Hehl^{§}

Institute for Theoretical Physics, University of Cologne

D-50923 Köln, Germany

Abstract

In order to be able to formulate a gravitational theory with nongravitating vacuum energy, Guendelman & Kaganovich introduced four scalar fields in spacetime. These fields are used to define a measure for integrating the Lagrangian 4-form over the spacetime manifold. We discuss the options one has for defining a volume element which can be used for physical theories. We show that one has to prescribe a scalar density σ. Whereas conventionally \sqrt{|det g_{ij}|} is used for that purpose, with g_{ij} as the components of the metric, we point out other possibilities, namely σ as a ‘dilaton’ field or as a derived quantity from either a linear connection or that quartet of scalar fields mentioned above.

∗Proceedings of the Workshop on Modern Modified Theories of Gravitation and Cosmology at Ben Gurion University, Beer Sheva, from 29-June to 30-June-1997. E.

Guendelman and A. Kaganovich (eds.), to be published in Hadronic Journal (1998).

†Email: fg@thp.uni-koeln.de ‡Email: muench@ph-cip.uni-koeln.de §Email: hehl@thp.uni-koeln.de

1 Introduction

The local geometry of spacetime is usually characterized by two independent concepts: The concept of a linear connection (parallel transport) and the concept of a metric (length and angle measurements). Both, the linear connection and the metric are physical fields which have to be determined from field equations of the gravitational theory under consideration.

In physics we do not only want to postulate the fields that are needed in order to formulate a theory. We also find it desirable to explain the existence of these fields by means of some fundamental principle. Therefore we are led to wonder what fundamental principles would suggest the existence of the linear connection and the metric, respectively.

The existence of the linear connection is quite satisfactorily explained by a symmetry principle which is believed to be fundamental: This is the principle of local gauge invariance. In the case of gravity we focus on external symmetries, i.e., symmetry transformations of spacetime. Local gauge invariance then reflects the invariance of a physical system under such transformations. This requires the introduction of a gauge connection which, in turn, allows to define parallel transport. The actual gauge connection depends on the specific symmetry under consideration and is, in a gauge approach to gravity, of the form of a linear connection together with the coframe; for details, see [10, 6].

In contrast to this, it is not clear how to derive the metric from some fundamental principle. Usually the existence of the metric is simply assumed, sometimes in disguise of a local symmetry group which contains an orthogonal subgroup. Therefore it is quite natural to ask whether the metric itself is a fundamental quantity, a derived quantity, or a quantity which can be substituted by some more fundamental field. To investigate this question is one of the motivations for this article. Another motivation is our desire to understand the work of Guendelman and Kaganovich (G&K for short) [7, 8, 9] and to put it into a proper perspective in view of the definition of the spacetime volume element. As we will see, both motivations will lead to interrelated questions and structures.

Our starting point is the observation that the metric is commonly taken to define a volume element in order to be able to perform integrations on spacetime. Integrated objects are clearly of fundamental importance in physics. However, the definition of a volume element on spacetime is also possible without reference to any metric. This general subject will be explained in Sec. 2. Basically, a volume element can be defined on any differentiable manifold as the determinant of a parallelepiped defined in terms of n vectors, if n is the dimension of the manifold. Then no absolute volume measure exists. However, proportions of different volumes can be determined. More explicitly, one finds that the volume element is the Levi-Civita n-form density transvected with the components of the n linearly independent vectors spanning the parallelepiped. Such a volume is an (odd) density of weight −1. In order to define an integral, we need then additionally a scalar density of weight +1.

Usual physical fields are no densities. Therefore the common practice is to take the metric and to build a density according to \sqrt{|detg_{ij}|}. But there exist alternatives, as we will point out in Sec. 3. They open the gate to alternative theories of gravitation.

Subsequently, in Sec. 4, we will follow up one possibility, namely the quartet of scalar fields, as proposed in [7, 8, 9]. In an appendix, we will provide some mathematical background for the differentiation of some quantities closely related to the volume element.

2 Integration on spacetime

We model spacetime as a 4-dimensional differentiable manifold, which is assumed to be paracompact, Hausdorff, and connected. We will restrict ourselves to four dimensions. Generalization to arbitrary dimensions is straightforward. In order to be able to formulate physical laws on such a spacetime, we have to come up with suitably defined integrals. If we want, for example, to specify a scalar action functional W of a physical system,

W = ∫ L = ∫ ϵ̃ L̂ ,  (1)

then, taking the integral in its conventional (Lebesgue) meaning, the Lagrangian L has to be an odd 4-form in order to make the integral (1) really a scalar. Incidentally, a p-form ω = \frac{1}{p!} ω_{i_{1}...i_{p}} dx^{i_{1}} ∧ ⋯ ∧ dx^{i_{p}} is called even if it is invariant under a diffeomorphism x^{i} → x′^{i}(x^{j}) with det(∂x^{j}/∂x′^{i}) < 0. It is called odd if it changes its sign under such a diffeomorphism; for even and odd forms, see Burke [2] and also [10].

Now, any odd 4-form can be split into a product of a 0-form (or scalar) L̂ and another odd 4-form ϵ̃. We assume that we did the splitting in such a way that the properties of the physical system are subsumed in the scalar L̂ and the properties of spacetime in the odd volume 4-form ϵ̃. In the case of gravity such a distinction may be ambiguous, because the properties of spacetime themselves are parts of the physical system. Let us consider a trivial physical system with L̂ = 1. Then the integral measures the volume of the corresponding piece of spacetime,

Vol = ∫ ϵ̃ .  (2)

For that reason ϵ̃ is called a volume form or, more colloquially, a volume element of spacetime. This quantity can be split again into two pieces. As the first piece we have the Levi-Civita ϵ in mind which is a very special geometric object. The Levi-Civita ϵ can be defined on any differential manifold. It is the ‘purely geometrical’ volume element. In order to define ϵ, we recall that, besides the components of the Kronecker symbol δ_{i}^{j}, the components ϵ_{ijkl} of the Levi-Civita ϵ are, by assumption, numerically invariant under diffeomorphisms. There exist no other quantities of this kind, apart from ϵ^{ijkl} of Sec. 3.4. And, in this sense, these components are very special. Now we define ϵ in terms of its components:

ϵ := \frac{1}{4!} ϵ_{ijkl} dx^{i} ∧ dx^{j} ∧ dx^{k} ∧ dx^{l} ,  with ϵ_{0123} = 1 = invariant . (3)

Consequently ϵ transforms as an odd 4-form density of weight −1 (see, e.g., [10, Appendix A], for details),

ϵ′ = \frac{1}{J} ϵ = \frac{sgnJ}{|J|} ϵ ,  (4)

where J = det(∂x^{j}/∂x′^{i}) is the determinant of the Jacobian matrix of the diffeomorphism x^{i} → x′^{i}(x^{j}). We are denoting densities by boldface letters.

We note in passing that for the manipulation of the ϵ-basis the following algebraic rules will turn out to be useful. If we take the interior product ⌋ of an arbitrary frame e_{α} with the Levi-Civita ϵ 4-form density, then we find a 3-form ϵ_{α}; if we contract again, we find a 2-form ϵ_{αβ}, etc.:

ϵ_{α} := e_{α}⌋ϵ = \frac{1}{3!} ϵ_{αβγδ} ϑ^{β} ∧ ϑ^{γ} ∧ ϑ^{δ} ,  (5a)
ϵ_{αβ} := e_{β}⌋ϵ_{α} = \frac{1}{2!} ϵ_{αβγδ} ϑ^{γ} ∧ ϑ^{δ} ,  (5b)
ϵ_{αβγ} := e_{γ}⌋ϵ_{αβ} = \frac{1}{1!} ϵ_{αβγδ} ϑ^{δ} ,  (5c)
ϵ_{αβγδ} = e_{δ}⌋ϵ_{αβγ} = e_{δ}⌋e_{γ}⌋e_{β}⌋e_{α}⌋ϵ .  (5d)

Here, the coframe ϑ^{β} is dual to the frame e_{α}, that is, e_{α}⌋ϑ^{β} = δ_{α}^{β}. The (ϵ , ϵ_{α} , ϵ_{αβ} , ϵ_{αβγ} , ϵ_{αβγδ}) represent a basis for the odd form densities of weight −1. It is called ϵ-basis and can be used to define a metric independent duality operation. Instead of lowering the rank of the ϵ’s, we can also increase their rank by exterior multiplication with the coframe ϑ^{μ} :

ϑ^{μ} ∧ ϵ_{α} = +δ_{α}^{μ} ϵ ,  (6a)
ϑ^{μ} ∧ ϵ_{αβ} = −δ_{α}^{μ} ϵ_{β} + δ_{β}^{μ} ϵ_{α} ,  (6b)
ϑ^{μ} ∧ ϵ_{αβγ} = +δ_{α}^{μ} ϵ_{βγ} − δ_{β}^{μ} ϵ_{αγ} + δ_{γ}^{μ} ϵ_{αβ} ,  (6c)
ϑ^{μ} ∧ ϵ_{αβγδ} = −δ_{α}^{μ} ϵ_{βγδ} + δ_{β}^{μ} ϵ_{αγδ} − δ_{γ}^{μ} ϵ_{αβδ} + δ_{δ}^{μ} ϵ_{αβγ} .  (6d)

For the ϵ̃ (which is an odd 4-form density of weight 0), formulae analogous to (5),(6) are valid. We have just to add twiddles to the ϵ’s. Since ϵ is an odd density of weight −1, we can split the volume element ϵ̃, if we postulate the existence of an even scalar density σ of weight +1, that is,

σ′ = |J| σ .  (7)

For our purpose here, we postulated an even scalar density, since the ϵ̃ in (1) and the Levi-Civita ϵ in (3) are both odd. Then eventually, equation (1) can be rewritten as

W = ∫ \underbrace{L}_{odd 4-f.} = ∫ \underbrace{ϵ̃}_{odd 4-f.} \underbrace{L̂}_{scalar}= ∫ \underbrace{ϵ}_{odd 4-f. density, weight −1} \overbrace{σ}^{even scalar density, weight +1} \underbrace{L̂}_{scalar} . (8)

The scalar density σ, in contrast to the Levi-Civita ϵ, must be specified in some way, before one can actually do physics on the spacetime manifold. It is here where gravity comes in.

3 Choices for the scalar density σ

3.1 Metric

It is conventional wisdom to choose the square root of the modulus of the metric determinant as the scalar density for building up the volume element:

0σ := \sqrt{|det g_{ij}|} .  (9)

As soon as a metric g = g_{ij} dx^{i} ⊗ dx^{j} is given – the gravitational potential of general relativity – we can define _{0}σ. In conventional integration theory, this is called the volume measure. We prefer to call it the metric volume measure and, accordingly, η := _{0}σ ϵ the metric volume element. Remember that Einstein, in his 1916 review paper of general relativity, see [11, p. 304], only admitted coordinates such that _{0}σ =^{∗} 1. This amounted to using the ‘purely geometrical’ volume element ϵ by constraining the free choice of the coordinates.

3.2 Dilaton field

Alternatively, we can promote the scalar density to a new fundamental field of nature, compare also the model developed in [10, Sec.6]. The value of such a density _{1}σ can be viewed as a scale factor of the volume element, see also [1, 16]. Thus, from a physical point of view, it is interesting to investigate the role of _{1}σ as a scaling parameter which realizes a scale transformation on a physical system.

Scale transformations as symmetry transformations play an important part in physics. In particular, in the context of cosmological models and the study of the unification of the four interactions, it is common to start from a scale invariant theory (for the sake of renormalizability, e.g.) and later derive the scales nowadays observed by some mechanism like spontaneous symmetry breaking. Clearly, scale invariance has to be carefully defined—and there are several possibilities. Well-known scale transformations are the Weyl transformations g_{ij} ⟶ e^{2Φ}g_{ij} which can be thought of as a rescaling of the metric. In conventional theories, where the volume element is taken as \sqrt{|det g_{ij}|}, a Weyl transformation scales the volume element in a definite way. This is important if Weyl-invariance of an action integral is considered. The possibility to take a metric independent scalar density _{1}σ in place of \sqrt{|det g_{ij}|} in order to build a proper volume element gives new opportunities to define a concept of scale invariance. In this context, the field _{1}σ is known as a dilaton field. For example, a replacement of \sqrt{|det g_{ij}|} by an independent _{1}σ will, a priori, ‘decouple’ the scaling properties of the volume element from the scaling properties of the (gravitational and matter) Lagrangian. Moreover, the gravitational field equations which are obtained from varying the metric, will, in general, be changed. In particular, the metrical energy momentum current will take a modified form, due to the absence of the factor \sqrt{|det g_{ij}|}. Here, the possibility of introducing the independent field _{1}σ can lead to a modification of this current and (possibly) related anomalies. In this sense, employing _{1}σ as independent scale parameter yields a lot of opportunities for physical creativity. We will come back to this question elsewhere.

3.3 Linear connection

In a pure connection ansatz, we prescribe a linear connection Γ_{α}^{β} = Γ_{iα}^{β}dx^{i} (but no metric!). Define, as usual, the curvature-2-form by

R_{α}^{β} = dΓ_{α}^{β} − Γ_{α}^{γ} ∧ Γ_{γ}^{β}  (10)

and the Ricci-1-form by

Ric_{α} := e_{β}⌋R_{α}^{β} = Ric_{iα} dx^{i} .  (11)

Then

2σ := \sqrt{|det Ric_{ij}|} ,  (12)

with Ric_{ij} = Ric_{iα} e_{j}^{α}, is a viable scalar density, as first suggested by Eddington [4, Sec. 92], compare also Schrödinger [15] and the more recent work of Kijowski and collaborators, see [3]. Likewise, the corresponding quantity based on the symmetric part of the Ricci tensor,

3σ := \sqrt{|det Ric_{(ij)}|} ,  (13)

also qualifies as a volume measure. Note that_{2}σ and_{3}σ may have singular points in such a theory as soon as the Ricci tensor or its symmetric part vanish. There seems to exist no criterion around which would prefer, say, 2σ, as compared to _{3}σ. Lately, such theories have been abandoned.

3.4 A quartet of scalar fields

In close analogy to the components ϵ_{ijkl} of the Levi-Civita ϵ, we can define the totally antisymmetric tensor density ϵ^{ijkl} of weight +1. We put its numerically invariant component ϵ^{0123} = −1.

Following G&K [7], we then can define

4σ := −ϵ^{ijkl} (∂_{i}φ^{(0)}) (∂_{j}φ^{(1)}) (∂_{k}φ^{(2)}) (∂_{l}φ^{(3)})
= −\frac{1}{4!} ϵ^{ijkl} ϵ_{ABCD} (∂_{i}φ^{A}) (∂_{j}φ^{B}) (∂_{k}φ^{C}) (∂_{l}φ^{D}) ,  (14)

where A, …, D are indices of interior space. This definition yields, for the volume 4-form

\overline{η} := _{4}σ ϵ ,  (15)

the following relations:

\overline{η} = dφ^{(0)} ∧ dφ^{(1)} ∧ dφ^{(2)} ∧ dφ^{(3)} = \frac{1}{4!} ϵ_{ABCD} dφ^{A} ∧ dφ^{B} ∧ dφ^{C} ∧ dφ^{D} . (16)

If we introduce the abbreviation

∂_{A} := \frac{∂}{∂φ^{A}} ,  (17)

then the duality of dφ^{A} and ∂_{B} can be expressed as follows:

dφ^{A} [∂_{B}] = δ_{B}^{A} .  (18)

In analogy to the set of equation (5), we define the 3-form and the 2-form

\overline{η}_{A} := ∂_{A}⌋\overline{η} , \overline{η}_{AB} := ∂_{B}⌋\overline{η}_{A} , etc.  (19)

Explicitly they read

\overline{η}_{A} = \frac{1}{3!}ϵ_{ABCD}dφ^{B} ∧ dφ^{C} ∧ dφ^{D} , \overline{η}_{AB} = \frac{1}{2!}ϵ_{ABCD}dφ^{C} ∧ dφ^{D} , etc. (20)

In analogy to (6) we have

dφ^{N} ∧ \overline{η}_{A} = +δ_{A}^{N} \overline{η} , dφ^{N} ∧ \overline{η}_{AB} = −δ_{A}^{N} \overline{η}_{B} + δ_{B}^{N} \overline{η}_{A} , etc. (21)

We contract (21) and find

\overline{η} = \frac{1}{4} dφ^{N} ∧ \overline{η}_{N} , \overline{η}_{A} = \frac{1}{3} dφ^{N} ∧ \overline{η}_{AN} , etc.  (22)

We differentiate (22):

d \overline{η} = −\frac{1}{4} dφ^{N} ∧ d \overline{η}_{N} , d \overline{η}_{A} = −\frac{1}{3} dφ^{N} ∧ d \overline{η}_{AN} , etc. (23)

Now, \overline{η}, as a 4-form, is closed:

d \overline{η} = 0 .  (24)

Provided dφ^{A} ≠ 0, we find successively,

d \overline{η}_{A} = 0 , d \overline{η}_{AB} = 0 , etc.  (25)

Using this information, we can partially integrate (22) and can prove that all these forms are not only closed, but also exact:

\overline{η} = d [\frac{1}{4} φ^{N} ∧ \overline{η}_{N}] , \overline{η}_{A} = d [\frac{1}{3} φ^{N} ∧ \overline{η}_{AN}] , etc.  (26)

Using (16) and (20), we find

\frac{∂ \overline{η}}{∂ dφ^{A}} = \overline{η}_{A} , \frac{∂ \overline{η}_{A}}{∂ dφ^{B}} = \overline{η}_{AB} , etc.  (27)

or, because of (26):

d \frac{∂ \overline{η}}{∂ dφ^{A}} = 0 , d \frac{∂ \overline{η}_{A}}{∂ dφ^{B}} = 0 , etc.  (28)

Since the corresponding “forces” vanish too, as can be seen from (16) and (20),

\frac{∂ \overline{η}}{∂ φ^{A}} = 0 , \frac{∂ \overline{η}_{A}}{∂ φ^{B}} = 0 , etc.,  (29)

we find an analogous result for the variational derivatives:

\frac{δ \overline{η}}{δ φ^{A}} = 0 , \frac{δ \overline{η}_{A}}{δ φ^{B}} = 0 , etc.  (30)

Similarly, we have

\frac{δ\overline{η}}{δϑ^{α}} = 0 , \frac{δ\overline{η}_{A}}{δϑ^{α}} = 0 , etc.  (31)

and

\frac{δ\overline{η}}{δg_{αβ}} = 0 , \frac{δ\overline{η}_{A}}{δg_{αβ}} = 0 , etc.  (32)

That the G&K-volume element is an exact form is the distinguishing feature of this ansatz. It is for that reason why the existence of a quartet of fundamental scalar fields is required instead of only one scalar field. Under these circumstances, the volume (2) can be expressed, via Stokes’ theorem, as a 3-dimensional surface integral which doesn’t contribute to the variation of the action functional.

4 The G&K-theory

Using (8) and the G&K volume element (16) and denoting the gravitational Lagrangian by V = V (g_{αβ}, ϑ^{α}, Q_{αβ}, T^{α}, R_{α}^{β}) and the matter Lagrangian by L_{m} = L_{m}(g_{αβ}, ϑ^{α}, Γ_{α}^{β}, Ψ, dΨ), the action W reads

W = ∫ (V + L_{m}) = ∫ \overline{η} \underbrace{(V̂ + L̂_{m})}_{scalar} = ∫ d χ (V̂ + L̂_{m}) .  (33)

Note that the 3-form χ, according to (26), explicitly reads χ := φ^{N} ∧\overline{η}_{N}/4. If we add a constant λ to the scalar Lagrangian, we find

∫ d χ (V̂ + L̂_{m} + λ) = W + λ ∫ d χ .  (34)

Since the 3-dimensional hypersurface integral ∫_{∂ Vol} χ doesn’t contribute to the variation, the scalar Lagrangian is invariant under the addition of a constant. This is the main tool of G&K in their model building. Variation with respect to φ^{A} yields the corresponding field equations

\frac{∂ (V + L_{m})}{∂φ^{A}} − d \frac{∂ (V + L_{m})}{∂ dφ^{A}} = 0 .  (35)

Suppose, as G&K in fact do, that V̂ and L̂_{m} do not depend on the quartet field at all,

\frac{∂ V̂}{∂ φ^{A}} = 0 , \frac{∂ V̂}{∂ dφ^{A}} = 0 ,  \frac{∂ L̂_{m}}{∂ φ^{A}} = 0 , \frac{∂ L̂_{m}}{∂ dφ^{A}} = 0 ,  (36)

then the field equations for the quartet field read

(V̂ + L̂_{m}) \frac{∂ \overline{η}}{∂φ^{A}} − d[(V̂ + L̂_{m}) \frac{∂ \overline{η}}{∂ dφ^{A}} ] = 0 .  (37)

The first term vanishes, since \overline{η} does not depend on φ^{A} explicitly, see (29). Then the Leibniz rule yields

d [(V̂ + L̂_{m}) \frac{∂ \overline{η}}{∂ dφ^{A}} ] = \frac{∂ \overline{η}}{∂ dφ^{A}} d (V̂ + L̂_{m}) + (V̂ + L̂_{m}) \underbrace{d\frac{∂ \overline{η}}{∂ dφ^{A}}}_{=^{(28)}0} = 0 .(38)

Provided φ^{A} ≠ 0 and dφ^{A} ≠ 0, we can conclude that

d (V̂ + L̂_{m}) = 0 ,  i.e., V̂ + L̂_{m} = const .  (39)

The gravitational field equations following from δg_{αβ} and δΓ_{α}^{β} are not disturbed by the existence of φ^{A}. Hence the usual metric-affine formalism applies in its conventional form (see [10], for recent developments cf. [13, 14, 12, 5]), but the field equation (39) for the scalar field quartet φ^{A} has to be appended. Perhaps surprisingly, it is only one equation since, in addition to (31) and (32), we trivially have

\frac{δ\overline{η}}{δΓ_{α}^{β}} = 0 , \frac{δ\overline{η}_{A}}{δΓ_{α}^{β}} = 0 , etc.  (40)

5 Conclusion

We can reproduce the essential features of the G&K-theory without the necessity to specify the gravitational first-order Lagrangian other than by the property (36).

Acknowledgments We are grateful to Eduardo Guendelman for inviting one of us to the workshop in Beer Sheva, and we would like to thank him and Alex Kaganovich for reading our manuscript and for useful remarks.

A Covariant exterior derivative of the ϵ-basis

For computations with the volume elements, it is convenient to introduce the differentials of the ϵ- and the ϵ̃-basis. As soon as a linear connection 1-form Γ_{α}^{β} = Γ_{iα}^{β}dx^{i} is given – there is no need of a metric for that – we find by covariant exterior differentiation of (5),

Dϵ_{α} = T^{μ} ∧ ϵ_{αμ} ,  (41a)
Dϵ_{αβ} = T^{μ} ∧ ϵ_{αβμ} ,  (41b)
Dϵ_{αβγ} = T^{μ} ∧ ϵ_{αβγμ} ,  (41c)
Dϵ_{αβγδ} = 0 .  (41d)

Here T^{α} := Dϑ^{α} is the torsion 2-form. For the σ–modified ϵ̃–basis, the computations run on the same track,

Dϵ̃_{α} = \frac{Dσ}{σ} ∧ ϵ̃_{α} + T^{μ} ∧ ϵ̃_{αμ} ,  (42a)
Dϵ̃_{αβ} = \frac{Dσ}{σ} ∧ ϵ̃_{αβ} + T^{μ} ∧ ϵ̃_{αβμ} ,  (42b)
Dϵ̃_{αβγ} = \frac{Dσ}{σ} ∧ ϵ̃_{αβγ} + T^{μ} ∧ ϵ̃_{αβγμ} ,  (42c)
Dϵ̃_{αβγδ} = \frac{Dσ}{σ} ∧ ϵ̃_{αβγδ} .  (42d)

the advantage being that this basis is composed of forms of weight 0, i.e., not of densities.

If a metric g is given additionally, then we can take _{0}σ as scalar density and find generally,

\frac{D_{0}σ}{_{0}σ} = −2 Q ,  (43)

with the Weyl covector Q := Q_{γ}^{γ}/4 and the nonmetricity 1-form Q_{αβ} := −Dg_{αβ}. It is then simple to rewrite (42) in terms of the metric volume element η := _{0}σ ϵ:

Dη_{α} = −2Q ∧ η_{α} + T^{μ} ∧ η_{αμ} ,  (44a)
Dη_{αβ} = −2Q ∧ η_{αβ} + T^{μ} ∧ η_{αβμ} ,  (44b)
Dη_{αβγ} = −2Q ∧ η_{αβγ} + T^{μ} ∧ η_{αβγμ} ,  (44c)
Dη_{αβγδ} = −2Q ∧ η_{αβγδ} .  (44d)

These equations turn out to be very helpful in conventional applications. However, in the G&K-theory, we have to forget (44) and to take recourse to _{4}σ.

Thus, analogously to (42), we have

D\overline{η}_{α} = \frac{D_{4}σ}{_{4}σ} ∧ \overline{η}_{α} + T^{μ} ∧ \overline{η}_{αμ} ,  (45a)

D\overline{η}_{αβ} = \frac{D_{4}σ}{_{4}σ} ∧ \overline{η}_{αβ} + T^{μ} ∧ \overline{η}_{αβμ} ,  (45b)

D\overline{η}_{αβγ} = \frac{D_{4}σ}{_{4}σ} ∧ \overline{η}_{αβγ} + T^{μ} ∧ \overline{η}_{αβγμ} ,  (45c)

D\overline{η}_{αβγδ} = \frac{D_{4}σ}{_{4}σ} ∧ \overline{η}_{αβγδ} .  (45d)

We find

\frac{D_{4}σ}{_{4}σ} = \frac{d_{4}σ}{_{4}σ} − Γ_{α}^{α} ,  (46)

but we were not able to derive a more compact expression for (46).

References

[1] W. Buchmüller and N Dragon: Einstein gravity from restricted coordinate invariance, Phys. Lett. B207 (1988) 292–294.

[2] W.L. Burke: Applied Differential Geometry (Cambridge University Press, Cambridge 1985).

[3] D. Catto, M. Francaviglia, J. Kijowski: A purely affine framework for unified field theories of gravitation, Bull. Acad. Polon. Sci. (Phys. Astron.) 28 (1980) 179–186.

[4] A.S. Eddington: The Mathematical Theory of Relativity, 2nd ed. (Cambridge University Press, Cambridge 1924).

[5] A. García, C. Lämmerzahl, A. Macías, E.W. Mielke, J. Socorro: Colliding waves in metric-affine gravity, Phys. Rev. D57 (1998), to be published. [gr-qc/9711041]

[6] F. Gronwald: Metric-affine gauge theory of gravity I. Fundamental structure and field equations, Int. J. of Mod. Phys. D6 (1997) 263–303. [gr-qc/9702034]

[7] E.I. Guendelman, A.B. Kaganovich: Principle of nongravitating vacuum energy and some of its consequences. Phys.Rev. D53 (1996) 7020–7025. [gr-qc/9605026]

[8] E.I. Guendelman, A.B. Kaganovich: Gravitational theory without the cosmological constant problem. Phys. Rev. D55 (1997) 5970–5980. [gr-qc/9611046] [9] E.I. Guendelman, A.B. Kaganovich: Gravitational theory without the cosmologi-

cal constant problem, symmetries of space filling branes and higher dimensions. Phys. Rev. D56 (1997) 3548–3554. [gr-qc/9702058]

[10] F. W. Hehl, J. D. McCrea, E. W. Mielke, Y. Ne’eman: Metric-affine gauge theory of gravity: Field equations, Noether identities, world spinors, and breaking of dilation invariance, Physics Reports 258 (1995) 1–171.

[11] A. J. Kox, M. J. Klein, R. Schulmann (editors): The Collected Papers of Albert Einstein, volume 6: The Berlin years: Writing 1914–1917 (Princeton University Press, Princeton 1996).

[12] A. Macías, E.W. Mielke, J. Socorro: Solitonic monopole solution in metric-affine gauge theory carrying Weyl charges, Class. Quantum Grav. 15 (1998), to be published.

[13] Y. Ne’eman, F.W. Hehl: Test matter in a spacetime with nonmetricity, Class. Quantum Grav. 14 (1997) A251–A260. [gr-qc/9604047]

[14] Yu.N. Obukhov, E.J. Vlachynsky, W. Esser, F.W. Hehl: Effective Einstein theory from metric-affine gravity models via irreducible decompositions, Phys. Rev. D56 (1997), to be published.

[15] E. Schrödinger: Space–Time Structure, reprinted with corrections (Cambridge University Press, Cambridge 1960).

[16] C. Wetterich: Cosmology and the fate of dilatation symmetry, Nucl. Phys. B302 (1988) 668–696.